# Great Models Think Alike and this Undermines AI Oversight

### v1

@@ B.3 Statistical Testing
B.3.1 Quantifying Correlation Strength Using Partial Correlation
@@ B.3.2 Multiple Regression
B.4 Experimental Setup for Filtering MMLU-Pro
We evaluate or models and judges on a set of questions that can be answered as MCQ as well as in open-style without access to reference options. This benchmark is obtained by using the filtering process proposed by Myrzakhan et al. (2024) on MMLU-Pro, whereas it was originally used to filter MMLU (Hendrycks et al., 2021; Wang et al., 2024). Every question is evaluated twice using a Qwen-2.5-32B-Instruct LM: first, it is judged in a binary way whether it is possible to answer the question without access to the MCQ options. In the second iteration, the judge gives a fine-grained confidence score. If either the binary decision is positive or the confidence is above a threshold, the question becomes part of our filtered benchmark. After this filtering process, 8707 of the original 12032 questions remain. The detailed prompts are described in Prompts B.8.5 and B.8.6.
B.5 Experimental Setup to Perform Free-Form Inference on Filtered MMLU-Pro
@@
We modified every 5-shot CoT prompt by removing the answer options from the end of the question and replacing every reference to them in the CoT with the corresponding answer text. An example of this process is shown in Prompts B.8.3 and B.8.4. Our benchmark is implemented as a task for the LM Eval Harness (Gao et al., 2023). Every CoT response is generated until a stop condition is met. The final response that is judged is extracted using regex matching. We use vLLM as the backend for the LM Eval Harness (Kwon et al., 2023). Even for instruction-tuned models, the options --apply_chat_template and --fewshot_as_multiturn were omitted because for the majority of LMs inspected the quality of responses decreased slightly to severely. However, we did not thoroughly investigate whether this is the case for every single model.
B.6 Experimental Setup for LLM-as-a-Judge on Filtered MMLU-Pro
@@
For the computation of similarities, we use the logs of the official evaluation runs of Myrzakhan et al. (2024) that are provided on huggingface.co. The set of responses is filtered to include only those questions that were rated as answerable in open-style without access to the reference options, as previously described in Section B.4. Using the logarithmic probabilities of the models for the answer options of this set of questions, we compute CAPA and other similarities. In addition, for each model-judge pair data samples where the ground truth option differed were excluded from the final analysis, for the final sample count per pair see from table 8 to table 16.
@@
B.7 List of Judges and Evaluated Language Models
@@
Whenever possible, we evaluated both the base and the instruction-tuned model for every combination of size and model family. Sometimes this was not possible, because the base model’s weights were not available on huggingface, evaluations on the Open LLM Leaderboard v2 were not provided or the LM consistently crashed in vLLM when performing inference (Myrzakhan et al., 2024; Kwon et al., 2023). The list below shows all models that are part of our experiments.
@@
Table 8: Final Sample Count (N) for Qwen2.5-7B-Instruct on Similarity Computation of Filtered MMLU-Pro
judge model N
Qwen2.5-7B-Instruct HuggingFaceTB/SmolLM2-1.7B 8707
HuggingFaceTB/SmolLM2-1.7B-Instruct 8706
HuggingFaceTB/SmolLM2-135M-Instruct 8707
HuggingFaceTB/SmolLM2-360M-Instruct 8707
Qwen/Qwen2.5-0.5B 8707
Qwen/Qwen2.5-0.5B-Instruct 8707
Qwen/Qwen2.5-1.5B 8707
Qwen/Qwen2.5-1.5B-Instruct 8707
Qwen/Qwen2.5-14B 8707
Qwen/Qwen2.5-14B-Instruct 8707
Qwen/Qwen2.5-32B 8707
Qwen/Qwen2.5-32B-Instruct 8707
Qwen/Qwen2.5-3B 8707
Qwen/Qwen2.5-3B-Instruct 8707
Qwen/Qwen2.5-72B 8707
Qwen/Qwen2.5-72B-Instruct 8707
Qwen/Qwen2.5-7B 8707
google/gemma-2-27b 8702
google/gemma-2-27b-it 8702
google/gemma-2-2b 8702
google/gemma-2-2b-it 8685
google/gemma-2-9b 8702
google/gemma-2-9b-it 8702
meta-llama/Llama-3.2-1B 8707
meta-llama/Llama-3.2-1B-Instruct 8707
meta-llama/Llama-3.2-3B 8707
meta-llama/Llama-3.2-3B-Instruct 8707
meta-llama/Llama-3.3-70B-Instruct 8706
meta-llama/Meta-Llama-3.1-70B 8685
meta-llama/Meta-Llama-3.1-70B-Instruct 8685
meta-llama/Meta-Llama-3.1-8B 8685
meta-llama/Meta-Llama-3.1-8B-Instruct 8702
microsoft/phi-4 8706
tiiuae/Falcon3-10B-Base 8706
tiiuae/Falcon3-10B-Instruct 8706
tiiuae/Falcon3-1B-Instruct 8706
tiiuae/Falcon3-7B-Base 8706
tiiuae/Falcon3-7B-Instruct 8706
Table 9: Final Sample Count (N) for Qwen2.5-32B-Instruct on Similarity Computation of Filtered MMLU-Pro
judge model N
Qwen2.5-32B-Instruct HuggingFaceTB/SmolLM2-1.7B 8707
HuggingFaceTB/SmolLM2-1.7B-Instruct 8706
HuggingFaceTB/SmolLM2-135M-Instruct 8707
HuggingFaceTB/SmolLM2-360M-Instruct 8707
Qwen/Qwen2.5-0.5B 8707
Qwen/Qwen2.5-0.5B-Instruct 8707
Qwen/Qwen2.5-1.5B 8707
Qwen/Qwen2.5-1.5B-Instruct 8707
Qwen/Qwen2.5-14B 8707
Qwen/Qwen2.5-14B-Instruct 8707
Qwen/Qwen2.5-32B 8707
Qwen/Qwen2.5-3B 8707
Qwen/Qwen2.5-3B-Instruct 8707
Qwen/Qwen2.5-72B 8707
Qwen/Qwen2.5-72B-Instruct 8707
Qwen/Qwen2.5-7B 8707
Qwen/Qwen2.5-7B-Instruct 8707
google/gemma-2-27b 8702
google/gemma-2-27b-it 8702
google/gemma-2-2b 8702
google/gemma-2-2b-it 8685
google/gemma-2-9b 8702
google/gemma-2-9b-it 8702
meta-llama/Llama-3.2-1B 8707
meta-llama/Llama-3.2-1B-Instruct 8707
meta-llama/Llama-3.2-3B 8707
meta-llama/Llama-3.2-3B-Instruct 8707
meta-llama/Llama-3.3-70B-Instruct 8706
meta-llama/Meta-Llama-3.1-70B 8685
meta-llama/Meta-Llama-3.1-70B-Instruct 8685
meta-llama/Meta-Llama-3.1-8B 8685
meta-llama/Meta-Llama-3.1-8B-Instruct 8702
microsoft/phi-4 8706
tiiuae/Falcon3-10B-Base 8706
tiiuae/Falcon3-10B-Instruct 8706
tiiuae/Falcon3-1B-Instruct 8706
tiiuae/Falcon3-7B-Base 8706
tiiuae/Falcon3-7B-Instruct 8706
Table 10: Final Sample Count (N) for Qwen2.5-72B-Instruct on Similarity Computation of Filtered MMLU-Pro
judge model N
Qwen2.5-72B-Instruct HuggingFaceTB/SmolLM2-1.7B 8707
HuggingFaceTB/SmolLM2-1.7B-Instruct 8706
HuggingFaceTB/SmolLM2-135M-Instruct 8707
HuggingFaceTB/SmolLM2-360M-Instruct 8707
Qwen/Qwen2.5-0.5B 8707
Qwen/Qwen2.5-0.5B-Instruct 8707
Qwen/Qwen2.5-1.5B 8707
Qwen/Qwen2.5-1.5B-Instruct 8707
Qwen/Qwen2.5-14B 8707
Qwen/Qwen2.5-14B-Instruct 8707
Qwen/Qwen2.5-32B 8707
Qwen/Qwen2.5-32B-Instruct 8707
Qwen/Qwen2.5-3B 8707
Qwen/Qwen2.5-3B-Instruct 8707
Qwen/Qwen2.5-72B 8707
Qwen/Qwen2.5-7B 8707
Qwen/Qwen2.5-7B-Instruct 8707
google/gemma-2-27b 8702
google/gemma-2-27b-it 8702
google/gemma-2-2b 8702
google/gemma-2-2b-it 8685
google/gemma-2-9b 8702
google/gemma-2-9b-it 8702
meta-llama/Llama-3.2-1B 8707
meta-llama/Llama-3.2-1B-Instruct 8707
meta-llama/Llama-3.2-3B 8707
meta-llama/Llama-3.2-3B-Instruct 8707
meta-llama/Llama-3.3-70B-Instruct 8706
meta-llama/Meta-Llama-3.1-70B 8685
meta-llama/Meta-Llama-3.1-70B-Instruct 8685
meta-llama/Meta-Llama-3.1-8B 8685
meta-llama/Meta-Llama-3.1-8B-Instruct 8702
microsoft/phi-4 8706
tiiuae/Falcon3-10B-Base 8706
tiiuae/Falcon3-10B-Instruct 8706
tiiuae/Falcon3-1B-Instruct 8706
tiiuae/Falcon3-7B-Base 8706
tiiuae/Falcon3-7B-Instruct 8706
Table 11: Final Sample Count (N) for Meta-Llama-3.1-8B-Instruct on Similarity Computation of Filtered MMLU-Pro
judge model N
Meta-Llama-3.1-8B-Instruct HuggingFaceTB/SmolLM2-1.7B 8702
HuggingFaceTB/SmolLM2-1.7B-Instruct 8701
HuggingFaceTB/SmolLM2-135M-Instruct 8702
HuggingFaceTB/SmolLM2-360M-Instruct 8702
Qwen/Qwen2.5-0.5B 8702
Qwen/Qwen2.5-0.5B-Instruct 8702
Qwen/Qwen2.5-1.5B 8702
Qwen/Qwen2.5-1.5B-Instruct 8702
Qwen/Qwen2.5-14B 8702
Qwen/Qwen2.5-14B-Instruct 8702
Qwen/Qwen2.5-32B 8702
Qwen/Qwen2.5-32B-Instruct 8702
Qwen/Qwen2.5-3B 8702
Qwen/Qwen2.5-3B-Instruct 8702
Qwen/Qwen2.5-72B 8702
Qwen/Qwen2.5-72B-Instruct 8702
Qwen/Qwen2.5-7B 8702
Qwen/Qwen2.5-7B-Instruct 8702
google/gemma-2-27b 8707
google/gemma-2-27b-it 8707
google/gemma-2-2b 8707
google/gemma-2-2b-it 8690
google/gemma-2-9b 8707
google/gemma-2-9b-it 8707
meta-llama/Llama-3.2-1B 8702
meta-llama/Llama-3.2-1B-Instruct 8702
meta-llama/Llama-3.2-3B 8702
meta-llama/Llama-3.2-3B-Instruct 8702
meta-llama/Llama-3.3-70B-Instruct 8701
meta-llama/Meta-Llama-3.1-70B 8690
meta-llama/Meta-Llama-3.1-70B-Instruct 8690
meta-llama/Meta-Llama-3.1-8B 8690
microsoft/phi-4 8701
tiiuae/Falcon3-10B-Base 8701
tiiuae/Falcon3-10B-Instruct 8701
tiiuae/Falcon3-1B-Instruct 8701
tiiuae/Falcon3-7B-Base 8701
tiiuae/Falcon3-7B-Instruct 8701
Table 12: Final Sample Count (N) for Meta-Llama-3.1-70B-Instruct on Similarity Computation of Filtered MMLU-Pro
judge model N
Meta-Llama-3.1-70B-Instruct HuggingFaceTB/SmolLM2-1.7B 8685
HuggingFaceTB/SmolLM2-1.7B-Instruct 8684
HuggingFaceTB/SmolLM2-135M-Instruct 8685
HuggingFaceTB/SmolLM2-360M-Instruct 8685
Qwen/Qwen2.5-0.5B 8685
Qwen/Qwen2.5-0.5B-Instruct 8685
Qwen/Qwen2.5-1.5B 8685
Qwen/Qwen2.5-1.5B-Instruct 8685
Qwen/Qwen2.5-14B 8685
Qwen/Qwen2.5-14B-Instruct 8685
Qwen/Qwen2.5-32B 8685
Qwen/Qwen2.5-32B-Instruct 8685
Qwen/Qwen2.5-3B 8685
Qwen/Qwen2.5-3B-Instruct 8685
Qwen/Qwen2.5-72B 8685
Qwen/Qwen2.5-72B-Instruct 8685
Qwen/Qwen2.5-7B 8685
Qwen/Qwen2.5-7B-Instruct 8685
google/gemma-2-27b 8690
google/gemma-2-27b-it 8690
google/gemma-2-2b 8690
google/gemma-2-2b-it 8707
google/gemma-2-9b 8690
google/gemma-2-9b-it 8690
meta-llama/Llama-3.2-1B 8685
meta-llama/Llama-3.2-1B-Instruct 8685
meta-llama/Llama-3.2-3B 8685
meta-llama/Llama-3.2-3B-Instruct 8685
meta-llama/Llama-3.3-70B-Instruct 8684
meta-llama/Meta-Llama-3.1-70B 8707
meta-llama/Meta-Llama-3.1-8B 8707
meta-llama/Meta-Llama-3.1-8B-Instruct 8690
microsoft/phi-4 8684
tiiuae/Falcon3-10B-Base 8684
tiiuae/Falcon3-10B-Instruct 8684
tiiuae/Falcon3-1B-Instruct 8684
tiiuae/Falcon3-7B-Base 8684
tiiuae/Falcon3-7B-Instruct 8684
Table 13: Final Sample Count (N) for Llama-3.3-70B-Instruct on Similarity Computation of Filtered MMLU-Pro
judge model N
Llama-3.3-70B-Instruct HuggingFaceTB/SmolLM2-1.7B 8706
HuggingFaceTB/SmolLM2-1.7B-Instruct 8707
HuggingFaceTB/SmolLM2-135M-Instruct 8706
HuggingFaceTB/SmolLM2-360M-Instruct 8706
Qwen/Qwen2.5-0.5B 8706
Qwen/Qwen2.5-0.5B-Instruct 8706
Qwen/Qwen2.5-1.5B 8706
Qwen/Qwen2.5-1.5B-Instruct 8706
Qwen/Qwen2.5-14B 8706
Qwen/Qwen2.5-14B-Instruct 8706
Qwen/Qwen2.5-32B 8706
Qwen/Qwen2.5-32B-Instruct 8706
Qwen/Qwen2.5-3B 8706
Qwen/Qwen2.5-3B-Instruct 8706
Qwen/Qwen2.5-72B 8706
Qwen/Qwen2.5-72B-Instruct 8706
Qwen/Qwen2.5-7B 8706
Qwen/Qwen2.5-7B-Instruct 8706
google/gemma-2-27b 8701
google/gemma-2-27b-it 8701
google/gemma-2-2b 8701
google/gemma-2-2b-it 8684
google/gemma-2-9b 8701
google/gemma-2-9b-it 8701
meta-llama/Llama-3.2-1B 8706
meta-llama/Llama-3.2-1B-Instruct 8706
meta-llama/Llama-3.2-3B 8706
meta-llama/Llama-3.2-3B-Instruct 8706
meta-llama/Meta-Llama-3.1-70B 8684
meta-llama/Meta-Llama-3.1-70B-Instruct 8684
meta-llama/Meta-Llama-3.1-8B 8684
meta-llama/Meta-Llama-3.1-8B-Instruct 8701
microsoft/phi-4 8707
tiiuae/Falcon3-10B-Base 8707
tiiuae/Falcon3-10B-Instruct 8707
tiiuae/Falcon3-1B-Instruct 8707
tiiuae/Falcon3-7B-Base 8707
tiiuae/Falcon3-7B-Instruct 8707
Table 14: Final Sample Count (N) for gemma-2-9b-it on Similarity Computation of Filtered MMLU-Pro
judge model N
gemma-2-9b-it HuggingFaceTB/SmolLM2-1.7B 8702
HuggingFaceTB/SmolLM2-1.7B-Instruct 8701
HuggingFaceTB/SmolLM2-135M-Instruct 8702
HuggingFaceTB/SmolLM2-360M-Instruct 8702
Qwen/Qwen2.5-0.5B 8702
Qwen/Qwen2.5-0.5B-Instruct 8702
Qwen/Qwen2.5-1.5B 8702
Qwen/Qwen2.5-1.5B-Instruct 8702
Qwen/Qwen2.5-14B 8702
Qwen/Qwen2.5-14B-Instruct 8702
Qwen/Qwen2.5-32B 8702
Qwen/Qwen2.5-32B-Instruct 8702
Qwen/Qwen2.5-3B 8702
Qwen/Qwen2.5-3B-Instruct 8702
Qwen/Qwen2.5-72B 8702
Qwen/Qwen2.5-72B-Instruct 8702
Qwen/Qwen2.5-7B 8702
Qwen/Qwen2.5-7B-Instruct 8702
google/gemma-2-27b 8707
google/gemma-2-27b-it 8707
google/gemma-2-2b 8707
google/gemma-2-2b-it 8690
google/gemma-2-9b 8707
meta-llama/Llama-3.2-1B 8702
meta-llama/Llama-3.2-1B-Instruct 8702
meta-llama/Llama-3.2-3B 8702
meta-llama/Llama-3.2-3B-Instruct 8702
meta-llama/Llama-3.3-70B-Instruct 8701
meta-llama/Meta-Llama-3.1-70B 8690
meta-llama/Meta-Llama-3.1-70B-Instruct 8690
meta-llama/Meta-Llama-3.1-8B 8690
meta-llama/Meta-Llama-3.1-8B-Instruct 8707
microsoft/phi-4 8701
tiiuae/Falcon3-10B-Base 8701
tiiuae/Falcon3-10B-Instruct 8701
tiiuae/Falcon3-1B-Instruct 8701
tiiuae/Falcon3-7B-Base 8701
tiiuae/Falcon3-7B-Instruct 8701
Table 15: Final Sample Count (N) for gemma-2-27b-it on Similarity Computation of Filtered MMLU-Pro
judge model N
gemma-2-27b-it HuggingFaceTB/SmolLM2-1.7B 8702
HuggingFaceTB/SmolLM2-1.7B-Instruct 8701
HuggingFaceTB/SmolLM2-135M-Instruct 8702
HuggingFaceTB/SmolLM2-360M-Instruct 8702
Qwen/Qwen2.5-0.5B 8702
Qwen/Qwen2.5-0.5B-Instruct 8702
Qwen/Qwen2.5-1.5B 8702
Qwen/Qwen2.5-1.5B-Instruct 8702
Qwen/Qwen2.5-14B 8702
Qwen/Qwen2.5-14B-Instruct 8702
Qwen/Qwen2.5-32B 8702
Qwen/Qwen2.5-32B-Instruct 8702
Qwen/Qwen2.5-3B 8702
Qwen/Qwen2.5-3B-Instruct 8702
Qwen/Qwen2.5-72B 8702
Qwen/Qwen2.5-72B-Instruct 8702
Qwen/Qwen2.5-7B 8702
Qwen/Qwen2.5-7B-Instruct 8702
google/gemma-2-27b 8707
google/gemma-2-2b 8707
google/gemma-2-2b-it 8690
google/gemma-2-9b 8707
google/gemma-2-9b-it 8707
meta-llama/Llama-3.2-1B 8702
meta-llama/Llama-3.2-1B-Instruct 8702
meta-llama/Llama-3.2-3B 8702
meta-llama/Llama-3.2-3B-Instruct 8702
meta-llama/Llama-3.3-70B-Instruct 8701
meta-llama/Meta-Llama-3.1-70B 8690
meta-llama/Meta-Llama-3.1-70B-Instruct 8690
meta-llama/Meta-Llama-3.1-8B 8690
meta-llama/Meta-Llama-3.1-8B-Instruct 8707
microsoft/phi-4 8701
tiiuae/Falcon3-10B-Base 8701
tiiuae/Falcon3-10B-Instruct 8701
tiiuae/Falcon3-1B-Instruct 8701
tiiuae/Falcon3-7B-Base 8701
tiiuae/Falcon3-7B-Instruct 8701
Table 16: Final Sample Count (N) for Ministral-8B-Instruct-2410 on Similarity Computation of Filtered MMLU-Pro
judge model N
Ministral-8B-Instruct-2410 HuggingFaceTB/SmolLM2-1.7B 8706
HuggingFaceTB/SmolLM2-1.7B-Instruct 8707
HuggingFaceTB/SmolLM2-135M-Instruct 8706
HuggingFaceTB/SmolLM2-360M-Instruct 8706
Qwen/Qwen2.5-0.5B 8706
Qwen/Qwen2.5-0.5B-Instruct 8706
Qwen/Qwen2.5-1.5B 8706
Qwen/Qwen2.5-1.5B-Instruct 8706
Qwen/Qwen2.5-14B 8706
Qwen/Qwen2.5-14B-Instruct 8706
Qwen/Qwen2.5-32B 8706
Qwen/Qwen2.5-32B-Instruct 8706
Qwen/Qwen2.5-3B 8706
Qwen/Qwen2.5-3B-Instruct 8706
Qwen/Qwen2.5-72B 8706
Qwen/Qwen2.5-72B-Instruct 8706
Qwen/Qwen2.5-7B 8706
Qwen/Qwen2.5-7B-Instruct 8706
google/gemma-2-27b 8701
google/gemma-2-27b-it 8701
google/gemma-2-2b 8701
google/gemma-2-2b-it 8684
google/gemma-2-9b 8701
google/gemma-2-9b-it 8701
meta-llama/Llama-3.2-1B 8706
meta-llama/Llama-3.2-1B-Instruct 8706
meta-llama/Llama-3.2-3B 8706
meta-llama/Llama-3.2-3B-Instruct 8706
meta-llama/Llama-3.3-70B-Instruct 8707
meta-llama/Meta-Llama-3.1-70B 8684
meta-llama/Meta-Llama-3.1-70B-Instruct 8684
meta-llama/Meta-Llama-3.1-8B 8684
meta-llama/Meta-Llama-3.1-8B-Instruct 8701
microsoft/phi-4 8707
tiiuae/Falcon3-10B-Base 8707
tiiuae/Falcon3-10B-Instruct 8707
tiiuae/Falcon3-1B-Instruct 8707
tiiuae/Falcon3-7B-Base 8707
tiiuae/Falcon3-7B-Instruct 8707
B.8 Prompts
B.8.1 LM-Judge Prompt without Reference Answer

### v2

@@ B.6 Statistical Testing
B.6.1 Quantifying Correlation Strength Using Partial Correlation
@@ B.6.2 Multiple Regression
B.7 Experimental Setup for Filtering MMLU-Pro
We evaluate or models and judges on a set of questions that can be answered as MCQ as well as in open-style without access to reference options. This benchmark is obtained by using the filtering process proposed by Myrzakhan et al. (2024) on MMLU-Pro, whereas it was originally used to filter MMLU (Hendrycks et al., 2021b; Wang et al., 2024). Every question is evaluated twice using a Qwen-2.5-32B-Instruct LM: first, it is judged in a binary way whether it is possible to answer the question without access to the MCQ options. In the second iteration, the judge gives a fine-grained confidence score. If either the binary decision is positive or the confidence is above a threshold, the question becomes part of our filtered benchmark. After this filtering process, 8707 of the original 12032 questions remain. The detailed prompts are described in Prompts B.11.5 and B.11.6.
B.8 Experimental Setup to Perform Free-Form Inference on Filtered MMLU-Pro
@@
We modified every 5-shot CoT prompt by removing the answer options from the end of the question and replacing every reference to them in the CoT with the corresponding answer text. An example of this process is shown in Prompts B.11.3 and B.11.4. Our benchmark is implemented as a task for the LM Eval Harness (Gao et al., 2023). Every CoT response is generated until a stop condition is met. The final response that is judged is extracted using regex matching. We use vLLM as the backend for the LM Eval Harness (Kwon et al., 2023). Even for instruction-tuned models, the options --apply_chat_template and --fewshot_as_multiturn were omitted because for the majority of LMs inspected the quality of responses decreased slightly to severely. However, we did not thoroughly investigate whether this is the case for every single model.
B.9 Experimental Setup for LLM-as-a-Judge on Filtered MMLU-Pro
@@
For the computation of similarities, we use the logs of the official evaluation runs of Myrzakhan et al. (2024) that are provided on huggingface.co. The set of responses is filtered to include only those questions that were rated as answerable in open-style without access to the reference options, as previously described in Section B.7. Using the logarithmic probabilities of the models for the answer options of this set of questions, we compute CAPA and other similarities. In addition, for each model-judge pair data samples where the ground truth option differed were excluded from the final analysis, for the final sample count per pair see from table LABEL:tab:Qwen2.5-7B-Instruct to table LABEL:tab:Ministral-8B-Instruct-2410.
@@
B.10 List of Judges and Evaluated Language Models
@@
Whenever possible, we evaluated both the base model and the instruction-tuned model for every combination of size and model family. Sometimes this was not possible, because the base model’s weights were not available on huggingface, evaluations on the Open LLM Leaderboard v2 were not provided or the LM consistently crashed in vLLM when performing inference (Myrzakhan et al., 2024; Kwon et al., 2023). The list below shows all models that are part of our experiments.
@@
B.11 Prompts
B.11.1 LM-Judge Prompt without Reference Answer